OpenStudy (anonymous):

Can someone please help me :( I don't know what to do I will become a fan, medal and everything. Please

OpenStudy (anonymous):

Here it is again

OpenStudy (kropot72):

(A) \[\frac{2}{3} \pi r _{1}^{3}=12 \pi r _{2}^{2}\] If you multiply both sides of the above equation by \[\frac{3}{2 \pi}\] the left hand side will become r1 cubed. Then you need to take the cube root of both sides, after first simplifying the right hand side.

OpenStudy (anonymous):

@kropot72 I'm still so confused. I don't get any of this. All of what you said flew over my head.

OpenStudy (anonymous):

@Hero ?

OpenStudy (kropot72):

@Don_n Do you want some help with algebra?

OpenStudy (anonymous):

@kropot72 yes

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (anonymous):

@imer ?